# Complex Physics-Informed Neural Network

### v1

@@ Abstract
We propose compleX-PINN, a novel physics-informed neural network (PINN) architecture that incorporates a learnable activation function inspired by Cauchy’s integral theorem. By learning the parameters of the activation function, compleX-PINN achieves high accuracy with just a single hidden layer. Empirical results show that compleX-PINN effectively solves problems where traditional PINNs struggle and consistently delivers significantly higher precision- often by an order of magnitude.
@@ 1 Introduction
Over the years, numerous strategies have been developed to enhance the performance of PINNs, including adaptive weighting of loss functions and selective sampling of training points. For example, Wang et al.[19] leveraged the Neural Tangent Kernel (NTK) to analyze gradient evolution in PINN training, adjusting the weights of each loss component accordingly. Other studies[21, 22] have explored methods for dynamically learning these weights during training. Additionally, adaptive sampling techniques have been introduced to tackle stiff problems by focusing on regions with high residuals. Lu et al.[23] proposed a threshold-based approach for selecting new training points, while Wu et al.[24] introduced a probability density function derived from residuals to improve sampling efficiency. Further extensions include high-dimensional adaptive methods [25] and re-sampling techniques targeting failure regions [26].
@@
Additional advancements include domain decomposition methods [29, 30, 31, 32], improved network initialization schemes [32, 33, 34], novel loss functions [35], and alternative network architectures [36, 37, 38, 39]. Collectively, these improvements have expanded the capabilities of PINNs for solving complex real-world problems.
However, existing methods primarily rely on additional training techniques or significantly larger models, such as transformers [36] and convolutional neural networks [39]. While these approaches can enhance performance, they also introduce substantial computational overhead. Auxiliary networks and gradient-based modifications [40, 41] increase training costs, whereas larger models suffer from slower convergence due to the large number of parameters [18, 42].
@@
The organization of this paper is as follows. Section 2 provides a brief introduction to PINNs. Our proposed model, compleX-PINN, is introduced in Section 3, where we first present the Cauchy activation function using Cauchy’s 1D integral formula in Section 3.1, extend it to high-dimensional cases in Section 3.2, and apply it to neural networks in Section 3.3. Finally, numerical results are presented in Section 4, followed by the conclusion.
@@ 3 Complex Physics-Informed Neural Network
where $\mu_{1}$, $\mu_{2}$, and $d$ are trainable parameters. This activation function is inspired by Cauchy’s integral formula, as we further elaborate in Section 3.1. We refer to a PINN model employing the Cauchy activation function as compleX-PINN.
We would like to note at the outset that our network is initially constructed with a single hidden layer, where each neuron has a unique set of parameters $\{\mu_{1},\mu_{2},d\}$. Consequently, the total number of trainable parameters for the Cauchy activation function is $3\times N_{\text{neuron}}$, where $N_{\text{neuron}}$ represents the width of the layer.
@@
Theorem 1 (Cauchy’s Integral Formula)
@@
At this point, we would like to highlight the power of the Cauchy approximation. For traditional activation functions such as ReLU or Sigmoid, the best one can hope for is a first-order approximation. However, with the Cauchy approximation, the error term is explicitly produced by approximating the contour integral with discrete values of the function. While we could have used higher-order Newton–Cotes formulas (such as Simpson’s rules or Boole’s rules) for numerical integration, they would yield exactly the same Cauchy approximation formula (9). Naturally, the values for $\lambda_{k}$ would vary slightly with different integration schemes, but since $\lambda_{k}$ are training variables meant to best fit the input data, this variation is inconsequential. Ultimately, the order of approximation can be arbitrarily high. In fact, for any integer $p$, the error term is estimated to be of the order $o(n^{-p})$.
@@ 3.2 Multi-dimensional Cauchy’s integral formula
Theorem 2 (Multi-Dimensional Cauchy’s Integral Formula)
@@
3.3 Extend the Cauchy activation function to high dimensional space
@@
where each $\Phi$ is a Cauchy activation function as defined in Equation (8). Here, the parameters $W_{k1},W_{k2},\ldots,W_{kN}$, $b_{k},\mu_{k1},\mu_{k2}$, and $d_{k}$ are trainable, allowing the network to capture the complex relationships among the input variables.
While the theoretical foundation of the approximation involves a single hidden layer, in practice, we can enhance the model’s capacity and expressiveness by constructing a neural network with multiple layers. This multilayer approach allows the model to capture hierarchical features and complex dependencies among the input variables, which are especially useful in high-dimensional cases.
Each hidden layer can be viewed as applying a set of nonlinear transformations to the input space, creating intermediate representations that capture interactions across multiple input variables. For a network with $L$ hidden layers, we recursively define each layer’s output as:
$h^{(l)}=\Phi\left(W^{(l)}h^{(l-1)}+b^{(l)};\mu^{(l)},d^{(l)}\right),$ (16)
where $h^{(l)}$ is the output of the $l$-th hidden layer, $W^{(l)}$ is the weight matrix of layer $l$, $b^{(l)}$ is the bias vector, $\mu^{(l)}$ and $d^{(l)}$ are the parameters of the Cauchy activation function for the $l$-th layer.
This recursive formulation allows the network to approximate increasingly complex functions as the number of layers grows, providing both depth and flexibility.
In this way, the network gains both practical advantages from multilayer structures and theoretical grounding from Cauchy’s integral approach, providing a novel architecture that can effectively approximate high-dimensional functions.
@@ 4 Numerical Experiments
In this section, we compare compleX-PINN with traditional PINN [1], Residual-Based Attention (RBA) PINN [42], and gradient-enhanced PINN (gPINN) [35] across various partial differential equations (PDEs), including the wave equation (Section 4.1), diffusion-reaction system (Section 4.2), Helmholtz equation (Section 4.3), reaction equation (Section 4.4), and convection equation (Section 4.5).
@@
CompleX-PINN is uniquely characterized by a single hidden layer, referred to as the Cauchy layer. Unlike other methods, each neuron in the Cauchy layer is parameterized by trainable parameters $\{\mu_{1},\mu_{2},d\}$, which provide greater flexibility in capturing complex solution patterns. Although each neuron introduces three additional parameters, the overall number of trainable parameters in compleX-PINN remains significantly smaller than multi-layer architectures used by competing methods, as a single hidden layer is sufficient for high performance. By default, the parameters $\{\mu_{1},\mu_{2},d\}$ are initialized to $0.1$, unless otherwise specified in particular experiments.
For testing, a uniform grid of points is generated for each dimension. For example, a $300\times 300$ grid is used for the 2D case, a $300\times 300\times 300$ grid for the 3D case, and this pattern is extended for higher-dimensional cases. Training points are then randomly sampled from these grids and are kept fixed across methods to ensure consistency. The number of training points is denoted as follows: $N_{f}$ for the residual, $N_{b}$ for the boundary conditions, and $N_{0}$ for the initial conditions.
To guarantee a fair comparison, all experiments are conducted with identical settings, including the learning rate, optimizer, and number of training epochs. Furthermore, the same random seed is used across all methods to enhance reproducibility. However, experimental setups, such as the number of training points or specific network architectures, may vary between PDEs to accommodate their unique characteristics.
Performance is evaluated using the relative $L^{2}$ error and the $L^{\infty}$ norm, which are defined as follows:
@@
where $u$ is the true solution of the PDE, $\hat{u}$ is the output from the tested method, and $N$ is the number of testing points.
4.1 Wave equation
We consider the following 1D wave equation:
$u_{tt}=4u_{xx},\quad$ $x\in[0,1],~{}t\in[0,1],$ (19)
$u(0,x)=\sin(\pi x)+\frac{1}{2}\sin(4\pi x),\quad$ $x\in[0,1],$ (20)
$u_{t}(0,x)=0,\quad$ $x\in[0,1],$ (21)
$u(t,-1)=u(t,1)=0,\quad$ $t\in[0,1],$ (22)
whose exact solution is given by:
$u(t,x)=\sin(\pi x)\cos(2\pi t)+\frac{1}{2}\sin(4\pi x)\cos(8\pi t).$ (23)
The wave equation is known to be challenging for PINNs due to its stiffness and the resulting difficulty in capturing high-frequency components of the solution [19, 21]. Traditional PINNs have slow convergence and suboptimal predictions. To address this challenge, the Residual-Based Attention (RBA) [42] has been proposed to improve PINNs by dynamically adjusting the weights of training points based on their residuals. RBA adaptively prioritizes regions of the solution domain where the model struggles the most, making it particularly effective for stiff PDEs. The weight $\lambda_{F_{i}}$ in Eq. (4) is updated iteratively using the following rule:
$\lambda_{F_{i}}^{k+1}=\gamma\lambda_{F_{i}}^{k}+\eta\frac{|r_{i}|}{\|r\|_{\infty}},$ (24)
where $\lambda_{F_{i}}$ is the weight associated with the training point $(\mathbf{x}_{i},t_{i})$, $r_{i}$ is the residual $r_{i}=\mathcal{F}[u](\mathbf{x}_{i},t_{i})$, and $r$ is the residual vector. The hyperparameters $\gamma$ and $\eta$ control the contribution of the residual term and its smoothing, and are set as in the original RBA implementation ($\eta=0.001$, $\gamma=0.999$). For additional details, refer to [42].
In this experiment, we set $N_{f}=6000$ and the boundary conditions and initial conditions are enforced by the hard constraint formulated as $\hat{u}=20x(1-x)t^{2}\hat{u}_{NN}+\sin{(\pi x)}+0.5\sin{(4\pi x)}$ [24, 46, 47], where $\hat{u}_{NN}$ represents the network output and $\hat{u}$ denotes the final prediction. For the standard PINN and RBA, we use a network with 5 hidden layers, each containing 300 neurons. For the compleX-PINN, we adopt a single hidden layer architecture with 1000 neurons. The number of parameters for PINN and RBA are $3\times 300+301\times 300\times 4+301\times 1=362,401$, while that for compleX-PINN is only $6\times 1000+1001\times 1=7,001$. Both models are trained using the Adam optimizer with a learning rate of $4\times 10^{-5}$ for 100k iterations.
The results of these experiments are presented in Fig. 1 and Fig. 2, with numerical results summarized in Table 1. As illustrated in Fig. 2, the standard PINN consistently exhibits a high relative $L^{2}$ error (more than 0.4) throughout training, highlighting its difficulty in solving stiff PDEs, as also reported in [19, 21]. When the Residual-Based Attention (RBA) technique is applied to the standard PINN, the relative $L^{2}$ error decreases rapidly, reaching 0.02 at 20k iterations and increasing to 0.08 at 100k iterations, showcasing the effectiveness of RBA in addressing these challenges.
In comparison, compleX-PINN alone achieves significantly more accurate predictions, reducing the relative $L^{2}$ error rapidly to below $0.01$ within just 20k iterations. When further enhanced with RBA (referred to as compleX-RBA), the relative $L^{2}$ error reduces from $3.01\times 10^{-3}$ to $1.14\times 10^{-3}$ after 100k iterations, demonstrating the synergistic effectiveness of combining compleX-PINN with RBA for solving stiff wave equations.
Figure 1: Prediction and pointwise error for the 1D wave equation using four methods: standard PINN (first row), RBA-PINN (second row), compleX-PINN (third row), and compleX-RBA (last row). CompleX-PINN achieves an $L^{\infty}$ error of $6.33\times 10^{-3}$, which is further reduced to $2.03\times 10^{-3}$ when the RBA technique is applied to compleX-PINN.
Table 1: Final performance of different models for the 1D wave equation after 100k iterations.
Model Relative $L^{2}$ error $L^{\infty}$ norm
PINN[1] $6.26\times 10^{-1}$ $7.56\times 10^{-1}$
RBA[42] $8.57\times 10^{-2}$ $2.30\times 10^{-1}$
compleX-PINN $3.01\times 10^{-3}$ $6.33\times 10^{-3}$
compleX-RBA $1.14\times 10^{-3}$ $2.03\times 10^{-3}$
Figure 2: Relative $L^{2}$ error history of PINN, RBA, compleX-PINN, and compleX-RBA for the 1D Wave equation. compleX-PINN outperforms both the standard PINN and RBA, and its performance is further improved when combined with the RBA technique.
4.2 Diffusion-reaction Equation
The diffusion-reaction equation, a parabolic PDE, models the macroscopic behavior of particles undergoing Brownian motion combined with chemical reactions. It finds applications in various fields, including information theory, materials science, and biophysics. In this section, we consider the following system, identical to the one presented in Ref. [35]:
$u_{t}=u_{xx}+R(x,t),\quad x\in[-\pi,\pi],\,t\in[0,1],$ (25)
$u(\pi,t)=u(-\pi,t)=0,$ (26)
$u(x,0)=\sum_{n=1}^{4}\frac{\sin(nx)}{n}+\frac{\sin(8x)}{8},$ (27)
where $R(x,t)$ represents the reaction term:
$R(x,t)=e^{-t}\left[\frac{3}{2}\sin(2x)+\frac{8}{3}\sin(3x)+\frac{15}{4}\sin(4x)+\frac{63}{8}\sin(8x)\right].$ (28)
The analytical solution to this system is given by:
$u(x,t)=e^{-t}\left(\sum_{n=1}^{4}\frac{\sin(nx)}{n}+\frac{\sin(8x)}{8}\right).$ (29)
For this experiment, we set $N_{f}=500$ and $N_{0}=N_{b}=50$. We compare compleX-PINN with RBA and Gradient-Enhanced PINN (gPINN) [35]. The compleX-PINN uses a hidden layer with 1500 neurons, while RBA and gPINN employ a 4-layer neural network with 100 neurons per hidden layer. All models are trained using the Adam optimizer with a learning rate of $1\times 10^{-5}$ for 15k iterations.
The predictions and pointwise errors are illustrated in Fig. 3, while the relative $L^{2}$ error histories are shown in Fig. 4. The results demonstrate that compleX-PINN converges much faster to a lower relative $L^{2}$ error compared to the other methods.
PINN: $2.05\times 10^{-1}(L^{\infty})$
gPINN: $9.30\times 10^{-2}(L^{\infty})$
RBA: $3.89\times 10^{-2}(L^{\infty})$
compleX-PINN: ${\mathbf{6.01\times 10^{-3}}}(L^{\infty})$
Figure 3: Pointwise error for the diffusion-reaction system using four methods: PINN, gPINN, RBA, and compleX-PINN. Among them, compleX-PINN exhibits the smallest $L^{\infty}$ error, while the errors of the other methods are more than six times larger.
Figure 4: Relative $L^{2}$ error history of different models for the diffusion-reaction equation. CompleX-PINN achieves a significantly lower relative $L^{2}$ error compared to all other models, and does so in fewer iterations. Specifically, compleX-PINN reduces the error below $1\times 10^{-2}$ within 5k iterations, whereas other models reach their lowest error of approximately $2\times 10^{-2}$ only after 15k iterations.
@@ 4.2 Diffusion-reaction Equation
The Helmholtz equation describes wave and diffusion phenomena, addressing variations over time within a spatial or combined spatial-temporal domain. We consider the following 2D Helmholtz equation

### v2

@@ Abstract
We propose compleX-PINN, a novel physics-informed neural network (PINN) architecture incorporating a learnable activation function inspired by Cauchy’s integral theorem. By optimizing the activation parameters, compleX-PINN achieves high accuracy with just a single hidden layer. Empirically, we demonstrate that compleX-PINN solves high-dimensional problems that pose significant challenges for PINNs. Our results show compleX-PINN consistently achieves substantially greater precision, often improving accuracy by an order of magnitude, on these complex tasks.
@@ 1 Introduction
Over the years, numerous strategies have been developed to enhance the performance of PINNs, including adaptive weighting of loss functions and selective sampling of training points. For example, Wang et al.[19] leveraged the Neural Tangent Kernel (NTK) to analyze gradient evolution in PINN training, adjusting the weights of each loss component accordingly. Other studies[21, 22] have explored methods for dynamically learning these weights during training. Additionally, adaptive sampling techniques have been introduced to tackle stiff problems by focusing on regions with high residuals. Lu et al.[23] proposed a threshold-based approach for selecting new training points, while Wu et al.[24] introduced a probability density function derived from residuals to improve sampling efficiency. Further extensions include active learning-based sampling methods [25] and re-sampling techniques targeting failure regions [26].
@@
Recent advancements in Physics-Informed Neural Networks (PINNs) include domain decomposition methods [29, 30, 31, 32], improved initialization schemes [32, 33, 34], novel loss functions [35], and innovative network architectures. For example, Kolmogorov-Arnold Networks (KANs) [36]—originally designed for general function representation—employ spline-based activation functions inspired by the Kolmogorov-Arnold theorem to enhance expressivity. Here, we adapt KANs for solving partial differential equations (PDEs), termed PIKAN [37, 38], to leverage their adaptive basis functions while addressing challenges in high-dimensional scalability. Alongside other architectural innovations like hierarchical networks [39, 40, 41, 42, 43], such advancements—including our PDE-focused adaptation of KANs—demonstrate the field’s progress in tailoring neural networks to the structure of physical systems. Collectively, these improvements have expanded PINNs’ utility for complex real-world applications, though optimizing expressivity and efficiency in high dimensions remains critical.
However, existing methods primarily rely on additional training techniques or significantly larger models, such as transformers [39] and convolutional neural networks [42]. While these approaches can enhance performance, they also introduce substantial computational overhead. Auxiliary networks and gradient-based modifications [44, 45] increase training costs, whereas larger models suffer from slower convergence due to the large number of parameters [18, 46]. Even recent innovations like KANs, despite their theoretical promise, face scalability challenges in high-dimensional PDE settings, as shown in our experiments (Section 4.4).
@@
The organization of this paper is as follows. Section 2 provides a brief introduction to PINNs. Our proposed model, compleX-PINN, is introduced in Section 3, where we first present the Cauchy activation function using Cauchy’s 1D integral formula in Section 3.1, extend it to high-dimensional cases in Section 3.2, and apply it to neural networks in Section 3.3. Finally, numerical results for testing several PDEs in high dimension are presented in Section 4, followed by the conclusion.
@@ 3 Complex Physics-Informed Neural Network
where $\mu_{1}$, $\mu_{2}$, and $d$ are trainable parameters. Moreover, $\mu_{1}$ governs the linear component of the input $x$, $\mu_{2}$ regulates the constant, enhancing adaptability in shaping the activation, and $d$ controls the activation function’s range and smoothness properties. The initialization effect of each parameters will be discussed later in Section 4.1.2. This activation function is inspired by Cauchy’s integral formula, as we further elaborate in Section 3.1. We refer to a PINN model employing the Cauchy activation function as compleX-PINN.
We would like to note at the outset that our network is initially constructed with a single hidden layer, where each neuron has a unique set of trainable parameters $\{\mu_{1},\mu_{2},d\}$. Consequently, the total number of trainable parameters for the Cauchy activation function is $3\times N_{\text{Cauchy}}$, where $N_{\text{Cauchy}}$ represents the width of the layer.
@@
Theorem 1 (Cauchy’s Integral Formula).
@@
At this stage, we emphasize the unique strength of the Cauchy activation function embedded in compleX-PINN. Unlike traditional activation functions, such as ReLU or Sigmoid, which typically offer only first-order approximation capabilities and are thus limited in expressiveness. The Cauchy activation function provides a natural path to high-order and even exponential approximation accuracy.
The key idea is that the activation function is derived from a discretized version of the Cauchy integral formula, where the network learns both the sample points and the associated weights. For analytic functions defined on compact domains, the approximation error of the Cauchy integral formula decays as $O(r^{-m})$ for some $r<1$, where $m$ is the number of sampling points and $r$ depends on the analyticity properties of the function and the geometry of its domain [50]. This exponential convergence is significantly faster than the algebraic decay $O(m^{-p})$ achieved by polynomial-based approximations of order $p$. In other words, the Cauchy activation function enables more efficient and accurate function representation, particularly for smooth or analytic functions.
While various numerical integration schemes—such as the trapezoidal rule, Simpson’s rule, Boole’s rule, or more general Newton–Cotes formulas—can be used to discretize the Cauchy integral, they ultimately produce the same functional form for the activation, up to the specific values of the quadrature weights $\lambda_{k}$ and nodes $z_{k}$. Since these quantities are treated as trainable parameters in the neural network, the choice of integration rule becomes inconsequential. The network adapts them to best approximate the target function during training.
@@ 3.2 Multi-dimensional Cauchy’s integral formula
Theorem 2 (Multi-Dimensional Cauchy’s Integral Formula).
@@
While the multi-dimensional Cauchy integral formula provides a powerful theoretical foundation for approximating holomorphic functions, its direct application in high-dimensional settings poses computational challenges. Specifically, the number of terms in the approximation grows exponentially with dimension $N$, and the resulting structure can be difficult to implement and optimize efficiently within standard neural network architectures. To address this, we propose a more practical alternative in the next section: approximating high-dimensional functions by applying the Cauchy activation function to suitable linear projections of the input. This approach not only retains the analytic expressiveness of the Cauchy activation function but also offers scalability and adaptability in learning complex features from high-dimensional data.
3.3 Using Cauchy activation to approximate high dimensional functions
@@
where each $\Phi$ is a Cauchy activation function as defined in Equation (8). Here, the parameters $W_{k1},W_{k2},\ldots,W_{kN}$, $b_{k},\mu_{k1},\mu_{k2}$, and $d_{k}$ are trainable, allowing the network to capture the complex relationships among the input variables. The validity of this approximation structure is grounded in the following theorem [47]:
Theorem 3 (General Approximation Theorem).
Let $\Phi$ be a family of functions in $\mathcal{C}(\mathbb{R},\mathbb{R})$, the space of continuous functions from $\mathbb{R}$ to $\mathbb{R}$, with the universal approximation property. Define
$\Phi^{N}=\left\{\phi_{a}(a_{1}x_{1}+a_{2}x_{2}+\cdots+a_{N}x_{N})\mid(a_{1},\ldots,a_{N})\in\mathbb{R}^{N},\phi_{a}\in\Phi\right\}.$
Then, the family $\Phi^{N}$ has the universal approximation property in $\mathcal{C}(\mathbb{R}^{N},\mathbb{R})$, i.e., every continuous function in $\mathbb{R}^{N}$ can be approximated by a linear combination of functions in $\Phi^{N}$, uniformly over compact subsets.
Since the Cauchy activation function $\Phi$ possesses the universal approximation property in one dimension (as shown in Equation (8)), Theorem 3 guarantees that the family in $\Phi^{N}$ (15) can approximate any continuous function in $\mathbb{R}^{N}$, uniformly over compact subsets. For completeness, we outline the proofs of this theorem based on Ref. [47].
Proof.
The set of all polynomials is dense in $\mathcal{C}(B,\mathbb{R})$ for any compact subset $B\subset\mathbb{R}^{N}$, by the Stone–Weierstrass Theorem. Since an analytical function $f$ on a compact domain can be approximated arbitrarily well by a polynomial, it suffices to show that each monomial can be approximated, up to an arbitrary order, by the Cauchy activation function.
As demonstrated earlier, the activation function $\Phi$ is capable of approximating one-dimensional monomials $f(x)=x^{k}$ for any integer $k$. This immediately implies that $\Phi$, when applied to linear combinations in $\mathbb{R}^{N}$, can approximate functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k},\;\mbox{ for }k\in\mathbb{N},\;(a_{1},\ldots,a_{N})\in\mathbb{R}^{N}.$
Fix an integer $k$, the multinomial expansion of the above function yields a linear combination of monomials:
$x_{1}^{k_{1}}x_{2}^{k_{2}}\cdots x_{N}^{k_{N}},\;\;k_{1}+k_{2}+\ldots+k_{N}=k.$
Conversely, by the Waring decomposition [51], any such multivariate monomial can itself be expressed as a linear combination of functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k}.$
Hence, the Cauchy activation function is capable of approximating any polynomial function and therefore any analytic function on compact domains in $\mathbb{R}^{N}$, establishing its expressive power in high dimensions. ∎
Cauchy activation offers a uniquely expressive framework for approximating complicated functions. A single layer with Cauchy activation achieves high precision in practice, often eliminating the need for multiple activation layers. While other architectural components (e.g., LSTMs) may still be necessary for domain-specific tasks [48], the simplicity of this design reduces computational overhead while maintaining accuracy.
@@ 4 Numerical Experiments
CompleX-PINN is uniquely characterized by a single hidden layer, referred to as the Cauchy layer. Unlike other neural networks, each neuron in the Cauchy layer is parameterized by trainable parameters $\{\mu_{1},\mu_{2},d\}$, which provide greater flexibility in capturing complex solution patterns. By default, the parameters $\{\mu_{1},\mu_{2},d\}$ are initialized to $0.1$, unless otherwise specified in particular experiments. Denote the number of neurons in the Cauchy layer as $N_{\text{Cauchy}}$, then the total number of parameters for a $D$-dimensional PDE with the 1-dimensional output is
$(D+1+3)\times N_{\text{Cauchy}}+(N_{\text{Cauchy}}+1)\times 1.$
Although each neuron introduces three additional parameters, the overall number of trainable parameters in compleX-PINN remains significantly smaller than currently used multi-layer architectures, as a single Cauchy layer is sufficient.
CompleX-PINN is evaluated against three state-of-the-art models representing distinct technical approaches:
  • Residual-Based Attention PINN (RBA-PINN) [46]. This method adaptively adjusts the residual loss weights at individual training points to emphasize more critical regions in the solution domain. It employs a standard fully connected neural network with depth $D_{\text{RBA}}$ and width $N_{\text{RBA}}$. The total number of trainable parameters is given by
    $(D+1)\times N_{\text{RBA}}+\sum_{i=1}^{D_{\text{RBA}}-1}(N_{\text{RBA}}+1)\times N_{\text{RBA}}+(N_{\text{RBA}}+1)\times 1.$
  • Binary structured PINN (BsPINN) [41]. This method employs a hierarchical binary architecture that systematically reduces inter-neuron connections, enhancing computational efficiency while maintaining expressive power. The network configuration is denoted as $M$-$\tilde{M}$, where the first hidden layer contains $M$ neurons grouped in a single block, and each subsequent layer halves the block size until it reaches $\tilde{M}$. This structure leads to $D_{\text{BsPINN}}:=\log_{2}(M/\tilde{M})+1$ hidden layers. The total number of parameters is calculated as
    $(D+1)\times M+\sum_{i=1}^{D_{\text{BsPINN}}-1}(2^{1-i}M+1)\times M+(M+1)\times 1.$
  • Physics-Informed Kolmogorov-Arnold Network (PIKAN) [36, 37, 38]. PIKAN [^1] The code of KAN is adapted from the efficientKAN https://github.com/Blealtan/efficient-kan. leverages the Kolmogorov-Arnold representation theorem by replacing traditional linear transformations with univariate B-spline functions. Given a spline order $p$ and $G$ intervals (which is equivalent to $G+1$ grid points), each univariate function is represented by $G+p$ basis functions. Each layer includes two additional learnable weight components—both initialized to $1$ [36]—corresponding to the base function and the spline function. These components contribute to the overall parameter count. With network width $N_{\text{KAN}}$ and depth $D_{\text{KAN}}$, as indicated in Ref. [38], the total number of parameters is
    $D\times N_{\text{KAN}}\times(G+p+2)+\sum_{i=1}^{D_{\text{KAN}}-1}N_{\text{KAN}}\times N_{\text{KAN}}\times(G+p+2)+N_{\text{KAN}}\times 1\times(G+p+2).$
The experimental setup is organized into distinct subsections, each targeting specific objectives, as outlined below:
  • Section 4.1: This section focuses on 2D Helmholtz problems, for which all three models have published results (albeit with varying PDE parameters). To ensure a fair and consistent comparison, we directly adopt the reported results from the original references, using identical training settings for each model. This avoids potential discrepancies due to reimplementation and preserves alignment with established baselines.
  • Section 4.2-4.4 (high-dimensional PDEs): In these sections, we explore high-dimensional PDEs by reproducing state-of-the-art methods using their publicly available codebases, adapted to our specific target equations. To ensure fair comparisons, we adhere to a standardized evaluation protocol across all methods. Moreover, to mitigate biases due to hyperparameter sensitivity and randomness in training, we perform extensive hyperparameter tuning for each model:
    • RBA-PINN: depth $D_{\text{RBA}}\in\{3,4,5\}$; width $N_{\text{RBA}}\in\{50,80,120,200\}$.
    • BsPINN: block size configurations: $\{128-8,128-16,256-16,256-32,512-32,512-64\}$.
    • PIKAN: number of intervals $G\in\{2,4,9\}$; width $N_{\text{KAN}}\in\{5,10,15\}$, with depth fixed as $D_{\text{KAN}}=2$ and spline order $p=3$, in accordance with common configurations used in prior works [37, 38]
    Following the tuning process, we report the experimental results corresponding to the best-performing configurations, selected based on validation performance.
Training and testing points. For 2D spatial or 1D spatio-temporal PDEs, test points are generated on a uniform $300\times 300$ grids (90,000 points). For higher-dimensional problems, 90,000 test points are randomly sampled to balance computational and memory constraints. Training points are also randomly selected, with $N_{f}$, $N_{b}$, and $N_{0}$ denoting the numbers of residual, boundary, and initial condition points, respectively. Note that training points are not necessarily included in the testing set.
For each PDE, each model is trained and evaluated across three independent trials using different random seeds. The best-performing result among the trials is reported. Performance is measured using the relative $L^{2}$ error and the $L^{\infty}$ norm, defined as follows:
@@
where $u$ denotes the ground truth solution, $\hat{u}$ is the prediction from the tested model, and $N$ is the number of testing points.
@@
The Helmholtz equation, widely used to model wave propagation and diffusion phenomena, describes physical processes that evolve over spatial or spatio-temporal domains. The two-dimensional (2D) Helmholtz equation has been extensively studied in the literature [46, 21, 22, 44, 52, 18, 37, 41]. Therefore, we compare compleX-PINN directly with existing results using the same problem setups reported in these studies.
We consider the following form of the 2D Helmholtz equation: